Two planes leave Dallas Airport at the same time, one flying east at a speed of 550 mph and the other flying west at the speed of 650 mph. How long does it take until they are 2000 miles apart?
#########
A helicopter leaves City Airport and flies north at 180 mph. Twenty minutes later a plane leaves the airport and follows the helicopter at 330 mph. How long does it take the plane to overtake the helicopter?
#########
A rectangle is 5 cm longer than it is wide. If its length and width are both increased by 3 cm, its area is increased by 60 cm squared. Find the dimensions of the original rectangle.

Answer by checkley77(12844)

550T+650T=2,000
1,200T=2,000
T=2,000/1,200
T=5/3 OR 1 HR. 40 MIN THEY WILL BE 2,000 MILES APART.
PROOF:
550*5/3+650*5/3=2,000
(2,750+3,250)/3=2,000
6,000=6,000
---------------------------------------
180T=330(T-1/3)
180T=330T-110
180T-330T=-110
-150T=-110
T=-110/-150
T=.7333 HOURS OR 44 MIN. TO OVERTAKE THE HELICOPTER AFTER IT LEAVES.
------------------------------------
L=W+5 ORIGINAL SIZE.
L+3 & W+3 IS THE EXPANDED SIZE.
LW=(L+3)(W+3)-60
(W+5)W=(W+3+5)(W+3)-60
W^2+5W=(W+8)(W+3)-60
W^2+5W=W^2+11W+24-60
5W-11W=24-60
-6W=-36
W=-36/-6
W=6 FOR THE ORIGINAL WIDTH.
L=6+5=11 FOR THE ORIGINAL LENGTH
11*6=66 ORIGINAL AREA.
(6+3)(11+3)=66+60
9*14=126
126=126